
At 02:54 we obtained the error: "Component(s) not ready for the next exposure: DECAL". We reset SISPI but did not work and Alberto went to the MZ floor to restat DECAL machine. After that, we pressed reset again in the Observer Console, the Interlock appeared green and then we go without any trouble. It took less than 5 minutes to do this. At 02:59 we were observing again.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2018-03-27 | 0.0 | 10.25 | 0.0 | 0.0 | 0.0 | 10.25 |
Links
[1] mailto:awalker@ctio.noao.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/5716/delete?destination=print/content/enr-ctio-blanco-4-m-2018-03-27
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/5716/edit?destination=print/content/enr-ctio-blanco-4-m-2018-03-27
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/13568?destination=print/content/enr-ctio-blanco-4-m-2018-03-27
SISPI was quite slow today. We have to kill and refresh the page from time to time.
I realized in the telemetry viewer that there was a peak at 04:30 in the LN2 Heater Power, Cryocoolers, Level, and LN2 diff pressure.